OpenStudy (anonymous):

(dz/dt) + 3e^(t+z) = 0 Medal for showing me this! :) I got z = -3e^t but it's wrong, can someone help please? This is for calc 2

OpenStudy (anonymous):

my assignment also asks to use C for any constant

OpenStudy (anonymous):

dz/dt=-3e^t*e^z e^-zdz=-3e^tdt integrating at both sides -e^-z=-3e^t+c e^-z=3e^t-c e^z=1/(3e^t-c)

OpenStudy (anonymous):

thank you I see what I did wrong now

OpenStudy (anonymous):

welcome!